# Layer 2 be or Layer not 2 be: Scaling on Uniswap v3

### 14 Mar 2024 v1

@@ 1 Introduction
Automated market makers (AMMs) - like the Uniswap Protocol - have become the dominant market structure for onchain markets, trading over \$3 trillion dollars to date. [^1] Source The Ethereum Protocol, currently the largest chain by onchain volume, has a limited availability for computation - or block space - and as user demand increases, the cost has risen due to limited supply. This is by design - block space needs to be constrained due to security and state growth. But as markets mature and the average transaction becomes more complex (increasing aggregate transaction costs), Ethereum’s base-layer throughput cannot feasibly keep up. As a result, there are comparatively fewer transactions on mainnet today than a few years ago, and those transactions are more expensive than they used to be.
@@
Lower costs will have significant benefits for Uniswap Protocol users. According to adams2023costs, 22% of the total cost of the average Ethereum transaction on the Uniswap Interface is Ethereum gas costs. On transactions less than \$1k, gas costs can make up as much as 98% of the total costs of trading. Lowering this cost allows users to pay less for the same service, potentially making trading costs comparable to centralized exchanges. Additionally, cheaper arbitrage transactions, and the related increase in fees paid by arbitrageurs, make LPs more profitable. It is also possible that users - both swappers and LPs - who are currently priced out of Ethereum mainnet would take part in the ecosystem if it were cheaper.
@@
Related Literature. Literature on automated market makers has grown rapidly in the past few years. Early roots can be traced back to hanson2007logarithmic and othman2013practical. Some recent literature about automated market makers include capponi2021adoption, lehar2021decentralized, and lvr2022. Uniswap v3 specifically has been described in Adams21 and empirically studied by adams2023costs. While there are relatively few papers on the empirics of Uniswap v3 outside of Ethereum mainnet, most likely due to the difficulty of generating the necessary data, there are a few notable contributions. chemaya2022cost presents findings showing that CPMMs on L2s give users better prices compared to Ethereum, but that users do not optimally switch based on the best total cost - showing a perceived security metric for the chain. caparros2023blockchain focuses on slippage and concentration between Polygon and Ethereum, showing similar benefits from downstream gas cost reductions. mitchell2022layer2 shows that the average full-range liquidity position does not have increased total fee returns on L2s, which will not contradict our results - as we focus only on arbitrage fee returns.
A technical discussion of L2s is out of the scope of this paper. For that, see thibault2022blockchain.
@@ 3 Benefits for Uniswap v3
There are two main categories of swappers on blockchain-based markets: arbitrageurs and retail flow. Aggregate statistics combine these two users, although they have different reasons for trading. Arbitrageurs will execute orders if they are profitable, and we will revisit their impact on market structure later. Retail flow is just as important (providing the bulk of positive returns for LPs in the $\mathsf{LVR}$ framework established in lvr2022) and are a bulk of the users who benefit from cheaper chains.
@@
The other cost of swapping for users is price impact, which is the movement of the “midmarket price” of the pool from execution of the trade. Uniswap v3 allows users to “concentrate liquidity” between prices, concentrating liquidity compared to a CPMM (like Uniswap v2). This concentration of liquidity has led to market depth dwarfing that of centralized exchange in some cases. For example, Uniswap v3 has two times the market depth of ETH/USD versus Binance - see liao2022dominance for more discussion on this.
@@
First, we will examine the difference in all-in swap cost of an USDC to ETH trade on Ethereum mainnet vs. Arbitrum. As previously stated, we will break down the cost into two factors: price impact and gas costs. We saw in Table 1 that gas costs on the studied L2s are much smaller, however overall liquidity is lower (and thus the price impact of a trade is higher). We utilize similar methodology to chemaya2022cost who calculate the optimal time for switching between chains.
@@
We have already shown the positive effects on market structure for LPs from lower gas - increased LP concentration, more unique liquidity providers, and decreased holding times. A second important aspect of the studied chains are their shorter block times. Historically, it has been difficult to calculate the benefit for market structure changes on accured LP fees, as LPs, arbitraguers, and swappers may respond differently to market structure changes. LPs are especially difficult due to utilizing different portfolio benchmarks. Liquidity providers may have arbitrary reasons for providing liquidity, and may benchmark against holding an arbitrary portfolio of tokens (impermanent loss), loss vs rebalancing ($\mathsf{LVR}$), or any other arbitrary strategies. However, we will focus on the well-researched $\mathsf{LVR}$, established in milionis2023automated, to compare between L1 and L2 LP performance. We also choose this because improving $\mathsf{LVR}$ generally improves returns on other strategies, and because the benchmark for the portfolio is the same between chains and at arbitrary times.
@@
From mitchell2022layer2, we can see there is very little difference between aggregate ETH/USDC fee returns of full-range positions on L1s and L2s. We know that there is more retail flow transacting on Ethereum, which may offset the decreased fee returns from arbitrage. We have also shown that retail traders continually trade on Ethereum mainnet - even though this is not optimal, likely because of lagged adoption of other chains. However, because of the coordination required between LPs and swappers, liquidity providers may stay on Ethereum because the fees from retail flow offset the increased costs from arbitrage (even if pareto optimal for both parties to be on L2s). Retail traders are the bulk of positive returns for liquidity providers, so liquidity providers must stay where they are.
@@
Next, we remove all swap transactions that originate with one of these interface routers (to remove retail flow). Lastly, we take all the remaining arbitrage transactions left, re-calculate their impact on state, and generate the full-range fee returns from all of these swaps. This is utilizes the feeGrowthGlobal values found in Uniswap v3 state. For proofs and instructions on replicating this methodology, see adams2022fees.
@@
As previously stated, the increase in fee-returns is from gas cost differences and shorter block times. However, during this sample period (done purposefully), Optimism and Arbitrum had nearly exact gas costs, while block times were the main difference. To zoom in on like-samples, we focused on March 1st, 2023 to December 1st, 2023. Over this period, we find that both Arbitrum and Optimism outperformed Ethereum by a significant margin - around 20% on average. Finally, by focusing only on arbitrage transactions (which will execute at any depth if profitable), we can remove the impact of liquidity differences of these pools between chains, meaning that equalizing the liquidity on L1s and L2s will not impact these dynamics. There is an interaction between full-range fee returns and increased concentration of other LPs, but it is minimal and due to gas costs. A deep dive into intra-pool LP dynamics is out of the scope of this paper - for that, see milionis2023flair.
Since September 2021, Ethereum mainnet returns for ETH/USDC would increase by about \$20m, surpassing some calculated losses to LPs during that time - meaning that Ethereum LPs on ETH/USDC could be in profit from only this change if they mark their returns with 5 minute $\mathsf{LVR}$ [^8] See here for query and here for data sources and canidio2023arbitrageurshere This assumes that Ethereum adopts the same block production mechanisms as L2s today (or that all capital and flow is perfectly migrated to L2s). For example, the FM-AMM from canidio2023arbitrageurs and associated CowSwap implementation on Ethereum mainnet claim increasing fees by 5-7% fixes $\mathsf{LVR}$, while simply using faster and cheaper chains increases it by an estimated 10-15%.
Importantly, the numbers for Ethereum include fee returns from sandwich transactions - which accounts for around 10% of ETH/USDC returns according to holloway2023value - meaning the difference is probably larger than 20%. Because sandwich bots utilize contracts which are generally also arbitrage bots, our methodology will include them as arbitrage transactions. Because of the current nature of sequencers, there are fewer sandwiches to no sandwiches on L2s, as sandwich trades without strict control of block ordering run the risk of being unbundled. [^9] For MEV number, see here and for a description on unbundling Accordingly, this removes sandwich transactions on the studied L2s, but also their associated LP returns.
@@ 4 Drawbacks for Uniswap v3
There is evidence that users internalize these implicit costs when interacting with L2s. chemaya2022cost estimated that L2 users internalize a 3.29% probability of total loss when trading CPMMs on Optimism and 0.68% on Polygon. They utilize the actual revealed preference of users switching between mainnet and the L2s to calculate this value. However, over time this number is trending towards 0, as the security tradeoffs of these systems lessen through the aforementioned improvements. The authors do note that other factors could impact this parameter, such as community, user experience, and overall integrations. We find similar results, as smaller users continue to execute trades on Ethereum mainnet despite the relatively better execution on L2s.
On the other hand, cong2023scaling finds L2s may lead to more secure markets, because fixed costs for interacting with markets are lower. Looking at the accuracy of Chainlink oracle feeds, Cong et al. estimate that costs for operators dropped by 76% while each round had 79% more operators, leading to a more secure, yet cheaper, feed. We find similar results with more unique liquidity providers providing liquidity on Uniswap v3 and smaller average Uniswap Interface trade size - signaling that users may be priced out of interacting with the mainnet ecosystems due to high gas costs.
@@ 5 Conclusion
Uniswap v3 was also optimized for Ethereum mainnet gas costs, meaning that the trading experience for the forthcoming Uniswap v4, described in Adams23, may benefit greatly from cheaper chains. Uniswap v4 allows for customization on top of the Protocol layer, which is up to the developer. Integrators on cheaper chains can utilize more complicated hooks, because of lower gas costs, to improve user experience. We have seen this already with hooks utilizing zero-knowledge proofs and Flashbots’ SUAVE, which were deployed on L2s and are likely not cost effective on mainnet.

### 15 Mar 2024 v2

@@ 1 Introduction
Automated market makers (AMMs) - like the Uniswap Protocol - have become the dominant market structure for onchain markets, trading over \$3 trillion dollars to date. [^1] Link to source The Ethereum Protocol, currently the largest chain by onchain volume, has a limited availability for computation - or block space - and as user demand increases, the cost has risen due to limited supply. This is by design - block space needs to be constrained due to security and state growth. But as markets mature and the average transaction becomes more complex (increasing aggregate transaction costs), Ethereum’s base-layer throughput cannot feasibly keep up. As a result, there are comparatively fewer transactions on mainnet today than a few years ago, and those transactions are more expensive than they used to be.
@@
Lower costs will have significant benefits for Uniswap Protocol users. According to Adams et al. (2023a), 22% of the total cost of the average Ethereum transaction on the Uniswap Interface is Ethereum gas costs. On transactions less than \$1k, gas costs can make up as much as 98% of the total costs of trading. Lowering this cost allows users to pay less for the same service, potentially making trading costs comparable to centralized exchanges. Additionally, cheaper arbitrage transactions, and the related increase in fees paid by arbitrageurs, make LPs more profitable. It is also possible that users - both swappers and LPs - who are currently priced out of Ethereum mainnet would take part in the ecosystem if it were cheaper.
@@
Related Literature. Literature on automated market makers has grown rapidly in the past few years. Early roots can be traced back to Hanson (2007) and Othman et al. (2013). Some recent literature about automated market makers include Capponi and Jia (2021), Lehar and Parlour (2021), and Milionis et al. (2022). Uniswap v3 specifically has been described in Adams et al. (2021) and empirically studied by Adams et al. (2023a). While there are relatively few papers on the empirics of Uniswap v3 outside of Ethereum mainnet, most likely due to the difficulty of generating the necessary data, there are a few notable contributions. Chemaya and Liu (2022) presents findings showing that CPMMs on L2s give users better prices compared to Ethereum, but that users do not optimally switch based on the best total cost - showing a perceived security metric for the chain. Caparros et al. (2023) focuses on slippage and concentration between Polygon and Ethereum, showing similar benefits from downstream gas cost reductions. Mitchell (2022) shows that the average full-range liquidity position does not have increased total fee returns on L2s, which will not contradict our results - as we focus only on arbitrage fee returns.
A technical discussion of L2s is out of the scope of this paper. For that, see Thibault et al. (2022).
@@ 3 Benefits for Uniswap v3
There are two main categories of swappers on blockchain-based markets: arbitrageurs and retail flow. Aggregate statistics combine these two users, although they have different reasons for trading. Arbitrageurs will execute orders if they are profitable, and we will revisit their impact on market structure later. Retail flow is just as important (providing the bulk of positive returns for LPs in the $\mathsf{LVR}$ framework established in Milionis et al. (2022)) and are a bulk of the users who benefit from cheaper chains.
@@
The other cost of swapping for users is price impact, which is the movement of the “midmarket price” of the pool from execution of the trade. Uniswap v3 allows users to “concentrate liquidity” between prices, concentrating liquidity compared to a CPMM (like Uniswap v2). This concentration of liquidity has led to market depth dwarfing that of centralized exchange in some cases. For example, Uniswap v3 has two times the market depth of ETH/USD versus Binance - see Liao and Robinson (2022) for more discussion on this.
@@
First, we will examine the difference in all-in swap cost of an USDC to ETH trade on Ethereum mainnet vs. Arbitrum. As previously stated, we will break down the cost into two factors: price impact and gas costs. We saw in Table 1 that gas costs on the studied L2s are much smaller, however overall liquidity is lower (and thus the price impact of a trade is higher). We utilize similar methodology to Chemaya and Liu (2022) who calculate the optimal time for switching between chains.
@@
We have already shown the positive effects on market structure for LPs from lower gas - increased LP concentration, more unique liquidity providers, and decreased holding times. A second important aspect of the studied chains are their shorter block times. Historically, it has been difficult to calculate the benefit for market structure changes on accured LP fees, as LPs, arbitraguers, and swappers may respond differently to market structure changes. LPs are especially difficult due to utilizing different portfolio benchmarks. Liquidity providers may have arbitrary reasons for providing liquidity, and may benchmark against holding an arbitrary portfolio of tokens (impermanent loss), loss vs rebalancing ($\mathsf{LVR}$), or any other arbitrary strategies. However, we will focus on the well-researched $\mathsf{LVR}$, established in Milionis et al. (2023a), to compare between L1 and L2 LP performance. We also choose this because improving $\mathsf{LVR}$ generally improves returns on other strategies, and because the benchmark for the portfolio is the same between chains and at arbitrary times.
@@
From Mitchell (2022), we can see there is very little difference between aggregate ETH/USDC fee returns of full-range positions on L1s and L2s. We know that there is more retail flow transacting on Ethereum, which may offset the decreased fee returns from arbitrage. We have also shown that retail traders continually trade on Ethereum mainnet - even though this is not optimal, likely because of lagged adoption of other chains. However, because of the coordination required between LPs and swappers, liquidity providers may stay on Ethereum because the fees from retail flow offset the increased costs from arbitrage (even if pareto optimal for both parties to be on L2s). Retail traders are the bulk of positive returns for liquidity providers, so liquidity providers must stay where they are.
@@
Next, we remove all swap transactions that originate with one of these interface routers (to remove retail flow). Lastly, we take all the remaining arbitrage transactions left, re-calculate their impact on state, and generate the full-range fee returns from all of these swaps. This is utilizes the feeGrowthGlobal values found in Uniswap v3 state. For proofs and instructions on replicating this methodology, see Adams and Liao (2022).
@@
As previously stated, the increase in fee-returns is from gas cost differences and shorter block times. However, during this sample period (done purposefully), Optimism and Arbitrum had nearly exact gas costs, while block times were the main difference. To zoom in on like-samples, we focused on March 1st, 2023 to December 1st, 2023. Over this period, we find that both Arbitrum and Optimism outperformed Ethereum by a significant margin - around 20% on average. Finally, by focusing only on arbitrage transactions (which will execute at any depth if profitable), we can remove the impact of liquidity differences of these pools between chains, meaning that equalizing the liquidity on L1s and L2s will not impact these dynamics. There is an interaction between full-range fee returns and increased concentration of other LPs, but it is minimal and due to gas costs. A deep dive into intra-pool LP dynamics is out of the scope of this paper - for that, see Milionis et al. (2023b).
Since September 2021, Ethereum mainnet returns for ETH/USDC would increase by about \$20m, surpassing some calculated losses to LPs during that time - meaning that Ethereum LPs on ETH/USDC could be in profit from only this change if they mark their returns with 5 minute $\mathsf{LVR}$ [^8] See here for query and here for data sources and Canidio and Fritsch (2023)here This assumes that Ethereum adopts the same block production mechanisms as L2s today (or that all capital and flow is perfectly migrated to L2s). For example, the FM-AMM from Canidio and Fritsch (2023) and associated CowSwap implementation on Ethereum mainnet claim increasing fees by 5-7% fixes $\mathsf{LVR}$, while simply using faster and cheaper chains increases it by an estimated 10-15%.
Importantly, the numbers for Ethereum include fee returns from sandwich transactions - which accounts for around 10% of ETH/USDC returns according to Holloway (2023) - meaning the difference is probably larger than 20%. Because sandwich bots utilize contracts which are generally also arbitrage bots, our methodology will include them as arbitrage transactions. Because of the current nature of sequencers, there are fewer sandwiches to no sandwiches on L2s, as sandwich trades without strict control of block ordering run the risk of being unbundled. [^9] For MEV number, see here and for a description on unbundling Accordingly, this removes sandwich transactions on the studied L2s, but also their associated LP returns.
@@ 4 Drawbacks for Uniswap v3
There is evidence that users internalize these implicit costs when interacting with L2s. Chemaya and Liu (2022) estimated that L2 users internalize a 3.29% probability of total loss when trading CPMMs on Optimism and 0.68% on Polygon. They utilize the actual revealed preference of users switching between mainnet and the L2s to calculate this value. However, over time this number is trending towards 0, as the security tradeoffs of these systems lessen through the aforementioned improvements. The authors do note that other factors could impact this parameter, such as community, user experience, and overall integrations. We find similar results, as smaller users continue to execute trades on Ethereum mainnet despite the relatively better execution on L2s.
On the other hand, Cong et al. (2023) finds L2s may lead to more secure markets, because fixed costs for interacting with markets are lower. Looking at the accuracy of Chainlink oracle feeds, Cong et al. estimate that costs for operators dropped by 76% while each round had 79% more operators, leading to a more secure, yet cheaper, feed. We find similar results with more unique liquidity providers providing liquidity on Uniswap v3 and smaller average Uniswap Interface trade size - signaling that users may be priced out of interacting with the mainnet ecosystems due to high gas costs.
@@ 5 Conclusion
Uniswap v3 was also optimized for Ethereum mainnet gas costs, meaning that the trading experience for the forthcoming Uniswap v4, described in Adams et al. (2023b), may benefit greatly from cheaper chains. Uniswap v4 allows for customization on top of the Protocol layer, which is up to the developer. Integrators on cheaper chains can utilize more complicated hooks, because of lower gas costs, to improve user experience. We have seen this already with hooks utilizing zero-knowledge proofs and Flashbots’ SUAVE, which were deployed on L2s and are likely not cost effective on mainnet.